The rms speed of the molecules in 1.5 g of hydrogen gas is 1700 m/s

What is the total translational kinetic energy of the gas molecules?

Express your answer using two significant figures.

What is the thermal energy of the gas?

Express your answer using two significant figures.

500 J of work are done to compress the gas while, in the same process, 1700 J of heat energy are transferred from the gas to the environment. Afterward, what is the rms speed of the molecules?

Express your answer using two significant figures.
Answer #1

  translational kinetic energy = 1/2mv^2 = 1/2(2.0g/1000g)(1700)^2 = 2900 j

Hydrogen isa diatomic gas, it has 5 degrees of freedom , so thermal energy= 5/3 kinetic energy

so thermal energy= 5/3 (2890) = 4800 j

(b) 4800 J(thermal energyinitial) + 500 J - 1700 = E thermalfinal = 5/3(1/2)mv2

3600 = (5/3)(1/2)(2/1000) V2

v= 1470 ms

so new rms speed (upto two sig figures) = 1500m/s